# Mitigate Position Bias in Large Language Models via Scaling a Single Dimension

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly applied in various real-world scenarios due to their excellent generalization capabilities and robust generative abilities. However, they exhibit position bias, also known as "lost in the middle", a phenomenon that is especially pronounced in long-context scenarios, which indicates the placement of the key information in different positions of a prompt can significantly affect accuracy. This paper first explores the micro-level manifestations of position bias, concluding that attention weights are a micro-level expression of position bias. It further identifies that, in addition to position embeddings, causal attention mask also contributes to position bias by creating position-specific hidden states. Based on these insights, we propose a method to mitigate position bias by scaling this positional hidden states. Experiments on the NaturalQuestions Multi-document QA, KV retrieval, LongBench and timeline reorder tasks, using various models including RoPE models, context window-extended models, and Alibi models, demonstrate the effectiveness and generalizability of our approach. Our method can improve performance by up to 15.2% by modifying just one dimension of hidden states.
@@ 1 Introduction
Long-context large language models (LLMs) (Gradient, 2024; Reid et al., 2024; Liu et al., 2024a; Young et al., 2024; Abdin et al., 2024; DeepSeek-AI, 2024) have recently garnered significant attention within the community, enabling LLMs to handle longer and more complex tasks such as long-context question-answering (Caciularu et al., 2023; Li et al., 2024b) and repository-level code understanding (Bairi et al., 2024). However, recent researches (Li et al., 2024b; Liu et al., 2024c; Li et al., 2024a; Shi et al., 2023; Tang et al., 2023), indicates that these long-context LLMs struggle to effectively and consistently utilize all the information provided in the context, exhibiting a position bias known as "lost in the middle", which means LLMs tend to ignore information in the middle of the prompt, even though they can utilize the information at the beginning and end of the prompts well. This issue occurs in nearly all LLMs (Liu et al., 2024c; Junqing et al., 2023; Zhang et al., 2024), whether they are decoder-only models or encoder-decoder models, powerful models or small LLMs. For example, for the GPT-3.5-Turbo model in the NaturalQuestion multi-document QA task, the performance difference between ground-truth information placed in the middle of the prompt versus at the ends is 22 points with 2.3k tokens prompt (Liu et al., 2024c). This significantly impacts the practical application of LLMs in real-world scenarios. Studies (Kamradt, 2023; Zhao et al., 2024) show that this position bias becomes more severe as the context length increases, hindering the practical application of long-context LLMs.
@@
Previous works have analyzed this issue from the perspectives of data distribution (Junqing et al., 2023; Yu, 2023; An et al., 2024) and position embeddings (Zhang et al., 2024; Chen et al., 2024). For example, FILM (An et al., 2024) addresses position bias by constructing data with key information distributed in various positions for supervised fine-tuning (SFT). Ms-PoE (Zhang et al., 2024) mitigates position bias by interpolating RoPE (Su et al., 2024) using head-wise scaling factors. However, these methods require additional overhead for training or online estimation of scaling coefficients and are currently applicable to only a few models, limiting their generalizability.
To fundamentally understand and alleviate position bias in LLMs, we first explored the micro-level manifestation of position bias in LLMs and observed patterns in the attention weights consistent with position bias. Next, we investigated the underlying causes of attention weight-induced position bias. By respectively modifying position embedding and causal mask, we found that, in addition to position embedding, the causal mask also significantly affects position bias. Further analysis revealed that the causal mask introduces "positional hidden states", which are positively correlated with absolute positions, thereby conveying positional information to LLMs. These positional hidden states appear regardless of what position encoding method is used, including RoPE (Su et al., 2024), Alibi (Press et al., 2022), and even NoPE (Haviv et al., 2022).
@@
Based on the above findings, we propose a position bias mitigation method named "scale positional hidden states". Specifically, we first design a prior-based searching algorithm that quickly identifies which dimensions of hidden states within the model are positional hidden states, using monotonicity, smoothness, and loss on validation sets as indicators. Next, we design an attention modification algorithm that only let the scaled hidden states influence the attention of the last token of the prompt, efficiently implemented using FlashAttention (Dao, 2024).
@@
Extensive experiments on various models, including LLaMA-2 (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023), Gemma (Team et al., 2024), Qwen (Bai et al., 2023), and MPT (Team, 2023), and across different tasks, including Multi-document QA, KV retrieval, LongBench (Bai et al., 2024) benchmark, and the timeline reorder task (Li et al., 2024a), demonstrate that our method effectively mitigates position bias by modifying only one dimension of the hidden states of the model, achieving improvements of up to 15.2%. Our method is compatible with various position embeddings, including RoPE (Su et al., 2024) and Alibi (Press et al., 2022), and shows good generalization.
@@
    We find that position bias can be reflected in attention patterns.
@@
    We discover that the causal mask also introduces position bias and generates positional hidden states correlated to absolute positions in the hidden layers.
@@
2 Beyond Position Embeddings: Positional Information can be seen in hidden states
In this section, we first identifies patterns in attention weights that closely correspond to position bias. Then, we discover that, apart from position embeddings, position information in the LLMs can also be generated by the causal mask, which tends to accumulate in a few specific hidden states channels and bears significant responsibility for the emergence of position bias.
@@ 2.1 Microscopic Manifestations of Position Bias in Transformers: Attention Weight Patterns
The attention of auto-regressive can be represented by the following equations:
$\begin{split}&\bm{q}=\mathcal{P}(W^{Q}\bm{h}(n),n),\quad\bm{k}=\mathcal{P}(W^{K}\bm{h}(m),m)\\ &\bm{a}_{n,m}=\text{Softmax}(\frac{\bm{q}\bm{k}^{\text{T}}+\text{Mask}}{\sqrt{d}})\end{split}$ (1)
where $\bm{h}$ is the hidden states, and $\bm{h}(n)$ is the hidden state of the n-th token. $W^{Q},W^{K}$ are the weights of the linear layers, $\mathcal{P}$ is the position encoding function like RoPE (Su et al., 2024), $d$ is the dimensionality of query and key states, and $n$ and $m$ are the positional order information. Mask is the causal mask.
To explore the micro-level manifestations of position bias in Transformers, we analyzed the attention weights for sentences containing key information, using a KV retrieval task, which requires the model to retrieval the ground-truth value of the given key from a list containing 50 Key-Value pairs (see Appendix B for details). As shown in Figures 1, in deep layers the model exhibits retrieval-like behavior, focusing on ground-truth information, forming a diagonal pattern observed in Figure 1b. While in other shallow layers, it always focus most attention on the start or end of the prompt, wherever the key information is located, exhibiting vertical lines patterns, as shown in Figure 1a.
In these layers exhibiting retrieval-like behavior, it can be observed that the attention weights for key information (Gold KV) exhibit patterns similar to position bias: when key information is located at the start or end of the prompt, the attention weights focused on it are relatively higher, while in the middle, they are significantly lower. Moreover, we extract the attention to key information (average of layers 15~25) with different context length in Figure 1c, where as the context length grows, the attenuation of attention weights with respect to position becomes more pronounced, reaching almost zero at the middle. More details about this are in Appendix E and B.
Furthermore, in Appendix 4.3, we found artificially adjusting the attention weights to the key information can directly improve the corresponding accuracy. Thus, we claim that position bias is to a large extent caused by the attention weights patterns at the micro level.
@@
Figure 1: Attention distribution of the ground-truth KV pair to each KV pair across different positions on the KV retrieval task (Liu et al., 2024c) using Mistral-7B (Jiang et al., 2023). (a) and (b) show the results averaged across all heads of the layer. (c) shows the attention of the ground-truth KV to the ground-truth KV (i.e., diagonal lines from (b)) across different context lengths.
@@
2.2 Causal Mask Also Contributes to Position Bias
@@
Figure 2: Performance of different methods with the ground-truth KV at different positions in the KV retrieval task (Liu et al., 2024c) using Mistral-7B (Jiang et al., 2023).
Based on Equ.(1), position embedding $\mathcal{P}$ allows LLMs to acquire postional information. However, recent works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) indicate that, besides position embeddings, the causal mask can also introduce positional information.
In this section, we aim to determine whether these two factors affect position bias by modifying different properties of the ground-truth KV pair. We introduce three baselines: (1) Crop Mask, which alters the causal mask so the ground-truth KV pair sees only itself, not previous tokens; (2) PE to Beginning, which assigns the position IDs of the ground-truth KV pair to match the first KV pair; (3) PE to End, which assigns the position IDs to match the last KV pair, details in Appendix C.
As shown in Figure 2, the original results exhibit a ”lost in the middle" pattern not only in accuracy but also in attention weight. Secondly, PE to end has a certain degree of help, but can hardly allow the model’s performance to match the accuracy when the ground-truth KV pair is positioned at the start or end of the prompt. Furthermore, PE to Beginning results in a noticeable performance drop as well as attention weight reduction when the gold KV is close to the end. In contrast, modifying the casual mask effectively enhances attention, especially to the latter KVs, and let the performance at the middle be improved to almost on par with the beginning. Based on the above observations, we can conclude that besides position embedding, the casual mask is also an important factor affecting position bias as well as corresponding attention weights. Moreover, solely modifying the position embedding hardly alleviates position bias completely.
2.3 Casual Mask Stores Position Information in Specific Hidden states Channels
Definition 2.1 (Positional Hidden States).
Let $h_{k}(p)$ denote the $k$-th dimension of the hidden states across each token’s position $p$. We define positional hidden states $h_{t}$ as hidden states whose values vary consistently and monotonically with the position sequence. Therefore, their derivative (after curve fitting) should always be positive or negative:
  • $h_{t}^{\prime}(p)>0$, $\forall p$ or $h_{t}^{\prime}(p)<0$, $\forall p$
Previous works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) have found that the positional information generated by causal mask is implicitly stored in hidden states. However, in fact, we find it can be observed explicitly, from “positional hidden states”.
To further analyze how positional information is transmitted in transformers, we define a special type of hidden state that directly reflects absolute positional information with high correlations to position IDs, called positional hidden states, as defined in Definition 2.1. We employ monotonicity rather than correlation as the primary property of positional hidden states, as correlation does not account for the sequential nature of positions. As shown in Figure 3, our experiments reveal that causal LLMs consistently possess such hidden states across most layers (details in Appendix F), even though these models do not have explicit absolute position embeddings, which means the causal mask is a very possible factor that provides absolute positional information. To demonstrate that these position hidden states are formed under the influence of the causal mask rather than the position embeddings, we conduct perturbation experiments on the causal mask and position embedding, as shown in Appendix C.
Based on the findings from Section 2.2, we conclude that the causal mask encodes positional information in certain hidden states, which subsequently influences attention weights and introduces position bias.
@@ 2.3 Casual Mask Stores Position Information in Specific Hidden states Channels
Figure 3: Averaged positional hidden states across all layers in different models.
3 Methodology
Based on the findings in Section 2, although the causal mask profoundly influences position bias, it is not feasible to know the positions of effective information in the prompt in advance, making methods that modify the causal mask difficult to design. Therefore, we propose a method to mitigate position bias by scaling the positional hidden states, as shown in Figure 4. Specifically, it consists of two steps: identifying the positional hidden states $h_{t}$ and scaling them by the factor $s$.
@@ 3.1 Problem Formulation
Given a pre-trained LLM $\bm{\theta}$ and a general dataset $\{\bm{x},\bm{y}\}$, our objective is to find the optimal positional hidden states $h_{t}$ and the corresponding scaling factor $s$ to maximally reduce position bias, which can be formulated as follows:
$\mathop{\arg\min}_{h_{t}\in\mathcal{H},s<1}\mathbb{E}\left[\sum_{i=1}^{|\bm{P}|}\mathcal{L}\left(\bm{x},\bm{y},\bm{p}_{i};F(\bm{\theta},h_{t},s)\right)\right]$ (2)
where $\bm{P}$ represents the set of different positions of the ground-truth information within the prompt $\bm{x}$, $F(\bm{\theta},h_{t},s)$ denotes the operation of scaling the LLM $\bm{\theta}$ on the $t$-th dimension of its hidden states by the scaling factor $s$, and $\mathcal{L}$ denotes the loss for general downstream tasks of the modified model.

### v3

# Mitigate Position Bias in LLMs via Scaling a Single Hidden States Channel

@@ Abstract
Long-context language models (LCLMs) can process long context, but still exhibit position bias, also known as “lost in the middle”, which indicates placing key information in the middle of the context will significantly affect performance. To mitigating this, we first explore the micro-level manifestations of position bias, concluding that attention weights are a micro-level expression of position bias. Then we identify that, in addition to position embeddings, positional information in hidden states also contributes to position bias, and it manifests itself in specific channels of hidden states, called positional hidden states. Based on these, we propose a method to mitigate position bias by scaling positional hidden states. Experiments on NaturalQuestions Multi-document QA, KV retrieval and LongBench, using various models including RoPE models, context window-extended models, and Alibi models, demonstrate the effectiveness and generalizability of our approach. Our method can improve performance by up to 15.2% in “lost in the middle” benchmark by modifying just one channel of hidden states. Our code is available at https://aka.ms/PositionalHidden.
Mitigate Position Bias in LLMs via Scaling a Single Hidden States Channel
Yijiong Yu1111Work during internship at Microsoft., Huiqiang Jiang2, Xufang Luo2, Qianhui Wu2, Chin-Yew Lin2, Dongsheng Li2, Yuqing Yang2, Yongfeng Huang1, Lili Qiu2 1Tsinghua University, 2Microsoft Corporation yuyj22@mails.tsinghua.edu.cn,yfhuang@tsinghua.edu.cn {hjiang,xufluo,qianhuiwu,cyl,dongsli,yuqyang,liliqiu}@microsoft.com
@@ 1 Introduction
Long-context language models (LCLMs) (Reid et al., 2024; Liu et al., 2025; Young et al., 2024; Abdin et al., 2024; DeepSeek-AI, 2024) have recently garnered significant attention within the community, enabling LLMs to handle longer and more complex tasks such as long-context question-answering (Caciularu et al., 2023; Li et al., 2025). However, recent research (Li et al., 2025; Liu et al., 2024; Li et al., 2024; Shi et al., 2023; Tang et al., 2023; He et al., 2024a; Zhang et al., 2024) shows that even long-context LLMs often fail to utilize all context information effectively, exhibiting a “lost in the middle" bias where middle context information is ignored. This issue affects all types of LLMs, regardless of their architecture or size, and worsens with longer contexts.
@@
Previous works have analyzed this issue from the perspectives of data distribution (He et al., 2024a; Yu, 2023; An et al., 2024) and position embeddings (Zhang et al., 2024; Chen et al., 2024). For example, FILM (An et al., 2024) addresses position bias by constructing data with key information distributed in various positions for supervised fine-tuning (SFT). Ms-PoE (Zhang et al., 2024) mitigates position bias by interpolating RoPE (Su et al., 2024) using head-wise scaling factors.
On the other hand, some works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) have proven that, besides position embeddings, the hidden states of LLMs can also convey positional information, which is generated by the causal attention mask. Although various works (see related works in Appendix A) have attempted to mitigate position bias, few of them have related position bias to positional information in hidden states, which may be another non-negligible source of position bias.
To verify the relation between position bias and positional information in hidden states, we conduct a series of experiments. First, as many previous works Yin et al. (2024); Wang et al. (2025); Chen et al. (2024) have observed, we find U-shaped attention weights consistent with position bias in specific layers. Second, we use perturbation experiments to prove position bias is indeed affected by positional information in hidden states. Third, We identify some channels called “positional channel” (the hidden states corresponding to them are called “positional hidden states”), whose values can manifest absolute positional information, through a process of hypothesis and subsequent verification. Finally, we confirm these channels can indeed affect position bias.
@@
Naturally, to mitigate position bias, a direct way is to eliminate (or minimize) its potential cause: positional hidden states. So, we propose a position bias mitigation method named “scale positional hidden states”. Specifically, we first design a heuristic searching algorithm that quickly identifies which channels (along the hidden size axis) of the hidden states are potential positional channel, using monotonicity and smoothness as indicators. Then We select the optimal channel among them based on the loss on a calibration dataset. Next, based on extensive engineering experience, we design a well-designed attention modification algorithm that let the scaled hidden states only influence the attention queried by the last token, to eliminate positional bias in attention, while avoiding interfering too much with the original attention leads to unstable model performance.
@@
Extensive experiments on various models, including LLaMA-2 (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Mistral (Jiang et al., 2023), Gemma (Team et al., 2024), Qwen (Bai et al., 2023), and MPT (Team, 2023), and across different tasks, including Multi-document QA, KV retrieval, LongBench (Bai et al., 2024) benchmark, demonstrate that our method effectively mitigates position bias by modifying only one channel of the hidden states of the model. In addition, we test on timeline reorder task (Li et al., 2024) and MMLU (Hendrycks et al., 2021) to show our methods has minimal side effects on the original capabilities of the model.
@@
    We are the first to discover that the positional information in hidden states is also the cause of position bias, especially the bias to the beginning.
@@
    We are the first to find explicit, visible hidden states channels which are approximately correlated to absolute token positions, called positional hidden states. And we confirm they can affect position bias.
@@
2 Positional Information in hidden states affects position bias
In this section, we first identify patterns in attention weights that closely correspond to position bias. Then, through modifying the causal mask in former layers and observing the change of attention in latter layers, we demonstrate position-related attention bias can also be affected by the positional information in hidden states. Third, we extract specific hidden states channels which has monotonicity with the token position, named “positional hidden states”, and find that they bears responsibility for the emergence of position bias. To sum up, we show our findings about position bias in this section in Figure 1.
Figure 1: The relationship between causal mask, positional information in hidden states, positional hidden states, position embedding, attention pattern and position bias of model’s performance.
@@ 2.1 Microscopic Manifestations of Position Bias in Transformers: Attention Weight Patterns
To explore the micro-level manifestations of position bias in Transformers, we analyze the attention weights in a classic long-context task, KV retrieval, which requires the model to retrieve the value of the given key from a list containing 50 Key-Value pairs (see Appendix D.1 for detailed prompts). The KV pair corresponding to the given key in the question is called the Gold KV. In attention analysis, we average all the attention weights from the last token of the question to the tokens of the $i$-th KV pair as the model’s “attention to the $i$-th KV”. More details about how we calculate attention are in Appendix D.1.
As shown in Figures 2, in deep layers, the model exhibits retrieval-like behavior, focusing on key information, forming a diagonal pattern observed in Figure 2b. So we call them retrieval-related layers. While in other shallow layers, it always focus most attention on the start or end of the prompt, wherever the key information is located, exhibiting vertical lines patterns, as shown in Figure 2a. Attention patterns of all the layers are shown in Appendix F, based on which we roughly regard that layers 15~20 belong to the retrieval-related layers.
In retrieval-related layers, it can be observed that the attention weights for key information (Gold KV) exhibit patterns similar to position bias: when key information is located at the start or end of the prompt, the attention weights focused on it are relatively higher, while in the middle, lower. Moreover, we extract the attention to key information (average of layers 15~20) with different context length. As shown in Figure 2c, as the context length grows, the attenuation of attention weights with respect to position becomes more pronounced.
Furthermore, in Appendix D.2, we find artificially adjusting the attention weights to the key information can directly improve the corresponding retrieval accuracy (this is intuitive). Thus, we confirm that position bias of model performance is to a large extent caused by the bias in attention weight patterns. So in the following experiments, we also use attention weights as the indicator of whether the model has successfully retrieved the key information, as well as the retrieval accuracy.
@@
Figure 2: Attention distribution of the gold KV pair to each KV pair across different positions on the KV retrieval task (Liu et al., 2024) using Mistral-7B (Jiang et al., 2023). (a) and (b) show the results averaged across all heads of the layer. (c) shows the attention of the ground-truth KV to the ground-truth KV (i.e., diagonal lines from (b)) across different context lengths.
@@
2.2 Positional Information in Hidden States Also Contributes to Position Bias
@@
Figure 3: Performance and attention of different methods with the ground-truth KV at different positions in the KV retrieval task (Liu et al., 2024) using Mistral-7B (Jiang et al., 2023).
Recent works (Haviv et al., 2022; Wang et al., 2024; Chi et al., 2023) have indicated that, besides position embeddings, the causal mask can also introduce positional information, which is then stored in the hidden states of LLMs. However, there is no research about whether such information will affect position bias. Therefore, we design a perturbation experiment using Mistral-7B-v0.2 (32 layers) on the classic KV retrieval task, where we modify the causal mask, aiming at changing position information in hidden states, and then observe whether position bias is affected.
In the experiment, only the causal mask of the layers 2 to 8 are modified, but we mainly observe the attention change in retrieval-related layers, specifically, the average attention of layers 15 to 20 is calculated. This setting avoids modifying the causal mask directly shifts the attention, letting us only see the effect of the positional information in hidden states.
The modification to the causal mask is called Crop Mask, which alters it so the tokens of the gold KV pair can only see itself, but not previous tokens (details about this operation are in Appendix D.3). Based on the theory of previous works, we posit this modification will make the positional feature of the tokens of the gold KV pair similar to that of the first KV pair.
In addition, as a comparison, we also shift the position embedding of the tokens of the gold KV (in every layer) as comparisons. Such shift includes: (1) PE to Beginning, which assigns the position IDs of the first KV pair to that of the gold KV pair; (2) PE to End, which assigns the position IDs of the last (the 50th) KV pair to it.
As shown in Figure 3, the original model apparently exhibits a “lost in the middle” pattern: when gold KV is in the middle position, the KV retrieval accuracy as well as the attention to it is much lower.
The most notable result is, modifying the casual mask effectively enhances the attention to the gold KV, as well as the retrieval accuracy, whatever its position is. It even lets the attention at the middle be improved to almost on par with the beginning.
As for PE modification, “PE to end” has a certain degree of improving attention to the gold KV, but can hardly allow the model’s performance to match the accuracy when the gold KV pair is positioned at the start or end of the prompt. In contrast, “PE to Beginning” results in a noticeable performance drop as well as attention weight reduction when the gold KV is originally close to the end.
This phenomenon indicates us that, the direct cause of the attention improvement in retrieval-related layers, can only be the information transmitted through hidden states (because PE and the causal mask in these layers are both not modified). That is, besides position embedding, the positional information in hidden states, which is introduced by causal mask, is also an important factor affecting position bias.
@@
Figure 4: Hidden states values with the token positions of the positional channel averaged across all layers.
2.3 Position Information can be clearly manifested in Specific Hidden states Channels
Definition 2.1 (Positional Hidden States).
Let $h_{k}(p)$ denote the $k$-th dimension of the hidden states across each token’s position $p$. We define positional hidden states $h_{t}$ as hidden states whose values change consistently and monotonically with the token position. Therefore, their derivative (after curve fitting) should always be positive or negative:
  • $h_{t}^{\prime}(p)>0$, $\forall p$ or $h_{t}^{\prime}(p)<0$, $\forall p$
We have found that the positional information in hidden states can affect position bias, and previous works Haviv et al. (2022) have proven positional information can be linearly probed from hidden states. However, the high-dimensional hidden states make it hard to concentrate on certain low-dimensional subspace to directly perceive or manipulate positional information in LLMs. To make this easier, we make a strong assumption: the positional information (at least part of it) in the hidden states should be manifested as activation values linearly correlated with each token’s absolute position. Thus, on average, there should exist some channels whose values change roughly monotonically with the absolute token positions, as shown in Definition 2.1. We call those channels positional channels and the parts of hidden states belonging to these channels as positional hidden states.
Based on this assumption, to identify positional channels, we average the model’s hidden states using 2000 randomly generated strings as inputs, whose lengths are 1000, to make an average hidden states with shape $[sequence\ length,hidden\ size]$. Then we traverse each channel (along $hidden\ size$ axis, in other words, $hidden\ size$ is the number of channels) to check whether its activation values (a one-dimensional array of $sequence\ length$) are changing monotonically with token positions. To eliminate the effect of outliers and noise, we apply sliding window average with window size of 100 to that array and discard the first 30 tokens (because the hidden states values of these tokens are often too huge (Sun et al., 2024)) before checking monotonicity.
As the results shown in Figure 4, even though our assumption is very strong, our experiments reveal that causal LMs consistently possess such hidden states across most layers (details in Appendix G), regardless of whether the model has position embeddings. We further demonstrate that position hidden states are mainly determined by the causal mask but not position embeddings through perturbation experiments in Appendix D.4.
Finally, we aim to confirm that even a single position channel can significantly affect position bias. So we conduct an experiment on KV retrieval, where we manually modify (subtract a fixed value from them) the values in the 213rd hidden states channel of Mistral-7b of a KV pair, and find the attention to this KV indeed greatly increases, which proves our posit. Details about this experiment are in Appendix D.5.
3 Mitigating Position Bias
Previous work focusing on PE often attempted to refine the application of RoPE, essentially minimizing the gap in position information (introduced by RoPE) between tokens in different positions.
In Section 2, we have identified positional information in hidden states as another factor causing position bias, and found positional channels strongly related to positional information. Thus, similarly, to minimize the gap in position information (introduced by positional information in hidden states) between tokens, the most direct way is to scale the activation values of those channels.
Therefore, we propose a method to mitigate position bias by scaling the positional hidden states, as shown in Figure 5. Specifically, it consists of two steps: identifying the positional hidden states $h_{t}$ and scaling (multiplying) them by a factor $s$.
@@ 3.1 Problem Formulation
Given a pre-trained LLM $\boldsymbol{\theta}$ and a general dataset $\{\bm{x},\bm{y}\}$, our objective is to find the optimal positional hidden states $h_{t}$, i.e. the $t$-th channel of the hidden states, and the corresponding scaling factor $s$ to maximally reduce position bias, which can be formulated as follows:
$\mathop{\arg\min}_{h_{t}\in\mathcal{H},s<1}\frac{1}{|\bm{P}|}\sum_{i=1}^{|\bm{P}|}\mathcal{L}\left(\bm{x},\bm{y},\bm{p}_{i};F(\boldsymbol{\theta},h_{t},s)\right)$ (1)
where $\bm{P}$ represents the set of different positions $p_{i}$ of the key information within the context of the prompt $\bm{x}$, $F(\boldsymbol{\theta},h_{t},s)$ denotes the operation of scaling the LLM $\boldsymbol{\theta}$ on the $t$-th channel ($t\in[1,hidden\ size]$) of its hidden states by the scaling factor $s$ (the specific scaling method is in Section 3.3), and $\mathcal{L}$ denotes the loss for general downstream tasks of the modified model.
In practice, due to the computing cost, the loss is computed on a small validation dataset of some representative tasks. Intuitively, we can traverse all different values of $t$ and obtain the loss on a calibration (validation) dataset to find the optimal one. However, due the large $hidden\ size$ which is typically over 4k, traversal search is too time-consuming, taking nearly 3 days for a 7B LLM. So, we design a search algorithm, mainly based on the monotonicity feature mentioned in Section 2, to first select a small set (no more than 50) of channels that are most likely to carry positional information that can affect position bias. Then we only need to traverse the small set of candidates to obtain their losses.